One charge of 2C is 1.5meters away from a -3C charge. Determine the force exerted on each other.

 

A -6 micro coulomb charge exerts an attractive force of 65N on a second charge that is 5cm away.What is the magnitude of the second charge?


1
Expert's answer
2022-02-20T15:58:41-0500

Solution;

(a)

"F=\frac{kq_1q_2}{r^2}"

"F=\frac{8.99×10^9×2×3}{1.5^2}=2.397×10^{10}N"

(b)

"65=\frac{8.99×10^9×6×10^{-6}×q_2}{0.05^2}"

"q_2=3.01\mu N"